OpenStudy (anonymous):

For breakfast, the cafeteria offers scrambled eggs with one slice of toast or an egg sandwich made with two slices of toast. In fifteen minutes, they sold 13 breakfasts and used 21 slices of bread. Use a model to determine how many egg sandwiches they sold. 6 7 8 9

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Let x = the number of scrambled egg plates sold and s = the number of egg sandwiches sold. Then x + s = 13 and x*1 + s*2 = 21 (where the number of each kind of breakfast is multiplied by the number of toast slices in each breakfast). Subtract the first equation from the second equation: x + 2s - x - s = 21 - 13 x - x + 2s - s = 8 s = 8

OpenStudy (anonymous):

Thank you

OpenStudy (doc.brown):

scram has 1 toast sam has 2 toast some scram and someother sam has 21 toast scram and sam has 13 breakfasts sub the scram and sam in the some scram and someother sam to get some 13 breakfasts minus sam and someother sam has 21 toast sam has 8 breakfasts